OpenStudy (anonymous):

If sin θ = 12/13 and cos ∅ = - 4/5, evaluate sin (θ-∅)

OpenStudy (apoorvk):

Okay okay.. You got \(\sin\theta\) and \(\cos\phi\), find out \(\cos\theta\) and \(\sin\phi\) using the identity \(\sin^2 x + \cos^2 x = 1\). Can you do this?

OpenStudy (diyadiya):

and.. \[\sin(x-y) =sinx~cosy-cosx~siny\]\[\sin(\theta - \phi ) = \sin \theta~\cos \phi - \cos \theta ~\sin\]

OpenStudy (diyadiya):

ooops

OpenStudy (apoorvk):

After this, use the identity: \[\sin(\theta - \phi) = \sin\theta\cos\phi - \cos\theta\sin\phi\]

OpenStudy (diyadiya):

\[\sin(x-y) =sinx~cosy-cosx~siny\]\[\sin(\theta - \phi ) = \sin \theta~\cos \phi - \cos \theta ~\sin \phi \]